OpenStudy (anonymous):

last one lol.... i need to go back and learn all this easier stuff i cant believe i forgot it all... Ruiz invests $55 at a simple interest rate of 4%. How much interest will he earn after 5 years

OpenStudy (anonymous):

Simple Interest Earned = P*R*T = 55 * 0.04 * 5 = 2.20 * 5 = $ 11